malik stops at the grocery store to buy a bag of diapers and 2 cans of formula. He spends a total of $37. The next week he stops and buys 2 bags of diapers and 5 cans of formula for a total of $87. How much does a bag of diapers cost? How much is one can of formula?

The sum of two numbers is -17. Their difference is 41. Find the numbers.
Answer #1

Let a bag of diapers is $ x and one can of formula is $ y .

Then ,

x +2y = 37 ...(1)

2x +5y = 87 ...(2)

(2) - (1)*2 implies

2x +5y -2x -4y =87 - 74

y = 13

Putting value of y in equation (1)

x + 2*13 = 37

x + 26 = 37

x = 37 -26

x = 11

A bag of diapers cost $ 11 and one can of formula is $ 13.